Grade 12th passAlgebra

Add:
3mn, – 5mn, 8mn, – 4mn
Add:
3mn, – 5mn, 8mn, – 4mn

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

When term have the same algebraic factors, they are like terms.
Then, we have to add the like terms
= 3mn + (-5mn) + 8mn + (- 4mn)
= 3mn–5mn + 8mn–4mn
= mn (3–5 + 8 - 4)
= mn (11 - 9)
= mn (2)
=2mn

Thanks